OpenStudy (anonymous):

WILL GIVE MEDAL A Web music store offers two versions of a popular song. The size of the standard version is 2.5 megabytes (MB). The size of the high-quality version is 4.8 MB. Yesterday, there were 1020 downloads of the song, for a total download size of 3378 MB. How many downloads of the standard version were there? Number of standard version downloads:

OpenStudy (sleepyjess):

Okay, well first we need to set up a system of equations. Do you know how to do that?

OpenStudy (anonymous):

No :(

OpenStudy (anonymous):

You want to make 2 equations with 2 variables. Lets call the amount of downloads for the standard version x, and the amount of download for the high quality version y. Now you know there is a total of 1020 download. This gives the first equation: \[x+y=1020\] Now you also know the size of the 2 versions, and the total download amount. This gives another equation: \[2.5x+4.8y=3378\] Now you need to solve the two equations with 2 variables. Do you know how to do that?

OpenStudy (anonymous):

@rfrisone

OpenStudy (anonymous):

I can solve the last equation

OpenStudy (anonymous):

What variable do we solve for @Tommynaiter

OpenStudy (anonymous):

You wan to solve both. So in order to find x and y, then we can do this by substitution. In this method you solve either x or y, in one of the equations and insert it in the next. \[x+y=1020 <=>x=1020-y\] Now we insert this on "x" in the 2nd equation: \[2.5*(1020-y)+4-8y=3378\] Now we have 1 equation with one variable. Can you solve y?

OpenStudy (anonymous):

If you want some more information/explanation about the method: https://en.wikipedia.org/wiki/System_of_linear_equations#Elementary_example